The number of proper subsets (123) is


 

Grade:

2 Answers

Arkalekha Neogi
11 years ago

The number of proper subsets is given by 2n-1,where n is the cardinal number of the set.

in A={1,2,3}, cardinal no. is 3 as number of elements is 3.

=>no. of proper subsets of A=2n-1 =23-1=8-1=7.

hope this is what you were looking for.

veeravalli mohan
11 years ago

the formula for number of proper sub sets is 2^n-2

so 

2^3-2

8-2=6

 

i hope you understood this